OpenStudy (anonymous):

What region R in the xy-plane maximizes the value of (integrand given below) Give reasons for your answer.

OpenStudy (anonymous):

\[\int\limits\limits_{}^{} \int\limits\limits_{R}^{} (4-x ^{2}-2y ^{2}) dA\]

OpenStudy (anonymous):

I know I should keep the postive integrand, but I don't know how to come up to the answer.

OpenStudy (dumbcow):

hmm if you set the function equal to 0, the result is an ellipse not sure if that works with double integrals though (maximize by differentiating and setting = to 0)

OpenStudy (anonymous):

@dumbcow How to differentiate?

OpenStudy (dumbcow):

thats what i mean the derivative of the integral is 4-x^2 -2y^2 but not sure if that applies for double integrals ?

OpenStudy (anonymous):

do you mean the answer is x^2 + y^2 <=4?